# Connecting Rashba and Dresselhaus spin-orbit interactions to inversion asymmetry in perovskite oxide heterostructures

### v1

@@ Abstract
Inversion asymmetry, combined with spin-orbit interaction, leads to Rashba or Dresselhaus effects or combinations that are promising for technologies based on antiferromagnetic spintronics. Since understanding the exact nature of spin-orbit interaction is crucial for developing a technology based on it, mapping the nature of inversion asymmetry with the type of spin-orbit interaction becomes the key. We simulate a perovskite oxide heterostructure LaAlOSrIrOSrTiO3 preserving the inversion symmetry within density functional theory to demonstrate the relation between the nature of inversion asymmetry and the corresponding Rashba or Dresselhaus-type interaction. With progressive distortion in the heterostructure, we find how the structure inversion asymmetry sets in with distorted bond lengths and angles, leading to the Rashba effect in the system. Further, the introduction of tilted IrO6 octahedra leads to bulk inversion asymmetry, helping a combined Rashba-Dresselhaus interaction to set in. A comparison of the spin textures obtained from our DFT calculations and theoretical modeling helps us identify the exact nature of the interactions. Besides demonstrating the connection between the nature of asymmetry and Rashba and Dresselhaus interactions, our work may serve as a guide to identifying different types of Rashba-like spin-orbit interactions.
@@ I Introduction
The surfaces and interfaces of perovskite oxides with the general formula ABO3 host Rashba-like interactions of various nature, order, and magnitude [10, 11, 12, 13, 14]. Additionally, owing to the two-dimensional conducting system at the interface and matured synthesis techniques, electronics using perovskite oxide heterostructures emerged as a promising research direction [15, 16]. Interesting magnetic properties have also been observed at such interfaces, shaping them as potential candidates for technology [17, 18, 11]. A host of possible distortions in perovskite oxides make them an ideal subject for studying several interesting physical properties and their interrelation with distortion. For example, KTaO3 is usually found in a perfect cubic structure, YMnO3 shows Jahn-Teller (JT) distortion, SrTiO3 exhibits tetragonal distortion, and CaTiO3 hosts tilted TiO6 octahedra. Additionally, inversion asymmetry arises in the structure due to the formation of interfaces or surfaces. Microscopic electric fields developed in the layered polar oxides also help manifest the Rashba interaction. Thus, perovskite oxide heterostructures offer a platform to systematically investigate and verify the specific types of asymmetry and their implications on Rashba-like spin-orbit interaction.
@@ II Method
The work has been carried out using ab initio calculations based on density functional theory (DFT) and analytical model calculations, with the latter being described in detail in the appendix A. The DFT calculations for the LaAlOSrIrOSrTiO3 heterostructures considered here were performed using the projector augmented wave (PAW) method [19] in conjunction with a plane-wave basis set, as implemented in the vasp code [20, 21]. A generalized gradient approximation (GGA) due to Perdew et al. [22] along with a Hubbard- correction for the transition metal -orbitals due to Dudarev et al. [23], henceforth referred to as the GGA+ method, was invoked for approximating the exchange-correlation functional, with a moderate value of , 3, and 10 eV for Ir-, Ti-, and La- orbitals, respectively [11]. Some key results were tallied against local density approximation (LDA) [24, 25] + Hubbard method, hereafter referred to as the LDA+ method. Unless mentioned otherwise for some specific cases, the lattice vectors and the atomic positions were optimized to minimize the stress and the Hellman-Feynman force on each atom to a tolerance of eV/Å, respectively, subject to a fixed lattice constant of 3.905 Å in the -plane to match that of a thick STO substrate in practice. -centered and -point meshes were used for the integrations over the Brillouin zone using improved tetrahedron method [26] with and cells, respectively, with switching to and -point meshes for the calculations of three-dimensional (3D) band dispersions and isoenergetic contours. Calculations accounting for spin-orbit interaction were performed with an energy convergence threshold of eV.
@@ III Results and Discussions
At first, we examine the spin-unpolarized density of states (DoS) and band structures of the heterostructure, shown in Fig. 2, in the three different structural forms mentioned above. The DoS and the projected DoS in Fig. 2(a) reveal the conducting nature of the heterostructure with the states near the Fermi level comprising Ir- orbitals hybridized with O- orbitals. The Ti- orbitals are found above the Fermi level, with the and orbitals higher in energy than the orbitals, as seen from the band dispersion in Fig. 2(b) [18, 10]. While the orbitals are degenerate in the bulk perovskites, confinement along the -direction in the heterostructure elevates the and bands in energy. From the electronic structure, we deduce that the nominal oxidation states for Ti and Ir are and , respectively. The structural form accommodating JT-like tetragonal distortion reveals no significant change in the electronics structure (see Fig. 2(c), Fig. 2(d)), except the and bands coming closer to the bands, owing to the elongation of bonds along the -direction, thereby reducing the effect of the confinement. The structure with optimally distorted bond lengths and bond angles exhibits some significant changes in the electronic structure, as seen from the DoS and the band dispersion in Fig. 2(e) and Fig. 2(f), respectively. We find the Ti- and the Ir- bands overlapping in energy above the Fermi level. The energy difference between the bands and the , bands further reduce upon accommodating the distortions, and some of the otherwise degenerate bands near the -point become non-degenerate due to reduced symmetry (see Fig. 2(f)).
@@
In order to understand the Rashba-like spin-orbit interaction in the system with progressive asymmetry, we plot the energy as functions of and (so-called 3D bands) corresponding to the bands crossing the Fermi energy for the heterostructures with all three types of distortions, and the corresponding isoenergetic contours and spin textures in Fig. 5. Fig. 5(a) displays the plot corresponding to the bands crossing the Fermi level for the undistorted heterostructure. Fig. 5(b) and Fig. 5(c) show the corresponding isoenergetic contours and the projected spin directions marked using arrows. All the 3D bands and the isoenergetic contours exhibited in Fig. 5 show a four-fold rotational symmetry, typical of and bands in the plane [10]. The projected spin vectors at a crystal momentum are almost collinear and do not show a helical pattern in this case. The essential characteristic of a spin texture corresponding to Rashba-like spin-orbit interaction, viz. for a particular band is also not observed here. Instead, the projected spins corresponding to a specific band exhibit behavior. Therefore, we conclude that the antiferromagnetic arrangement of spins leads to the opposite spins from different bands in the spin texture, while no Rashba-like interaction is observed in the undistorted heterostructure due to the symmetry being preserved. Owing to preserved overall inversion symmetry and centrosymmetry at the point group level, the heterostructure with JT-like tetragonal distortions also does not exhibit any Rashba-like interaction, as confirmed from the dispersion plot, the isoenergetic contours, and the spin textures shown in Fig. 5(d),(e),(f). The Rashba-like spin helix obeying the condition for a particular band is not observed here, although the spin vectors far away from the Brillouin zone center are not always collinear with those near the zone center.
In the case of the heterostructure with distorted bond lengths and bond angles, we find from the 3D bands in Fig. 5(g), and the isoenergetic contours in Fig. 5(h) and Fig. 5(i) that the bands crossing the Fermi level are almost degenerate at and lines, but have a large separation along the diagonals of the Brillouin zone. The projected spin textures corresponding to these bands, displayed in Fig. 5(h) and Fig. 5(i), reveal the spins orienting themselves tangentially to a circle, obeying the feature. Such orientation of the spins confirms predominantly linear Rashba-like interaction in the system (see the appendix) owing to the compromised structure inversion symmetry due to the distortions.
@@ IV Conclusion
Our study illustrates how inversion asymmetry can lead to Rashba and Dresselhaus spin-orbit interaction using an example of a perovskite oxide heterostructure. Our first-principles calculations within density functional theory, 3D band dispersion, and a comparison of the spin textures obtained from our DFT calculations and theoretical models help us connect the structural distortions with the Rashba-like spin-orbit interaction pattern observed in such systems. Starting from a heterostructure of LaAlOSrIrOSrTiO3 simulated in such a way that preserves the inversion symmetry, we allowed progressive distortions in the bond lengths and the bond angles, without the IrO6 octahedra being tilted. Our results reveal linear Rashba spin-orbit interaction for the distorted heterostructure, suggesting only structure inversion asymmetry was realized through the distortions, while the bulk inversion symmetry was intact. However, considering a heterostructure allowed for tilted IrO6 octahedra, leading to bulk and structure inversion asymmetry manifesting as a combined Rashba and Dresselhaus spin-orbit interaction. A comparison of the spin texture obtained from our DFT calculations for a heterostructure and a spin texture predicted from our previous model calculation are in excellent agreement [11], implying the model’s validation. We have also considered various combinations of Rashba and Dresselhaus interactions, compiling an elaborate reference for the spin-splitting and spin textures corresponding to such interactions. Besides illustrating the connection between structural asymmetry and the Rashba-like spin-orbit interactions through an example of perovskite oxide heterostructures, our work helps classify via the spin textures similar spin-splitting due to spin-orbit interaction. Our techniques of analyzing the spin textures may help study topological materials and understand the corresponding transport properties.
@@ Appendix A A manual for identifying the nature of Rashba-like interactions
Various Rashba-like spin-orbit interactions of different orders and their combinations may be identified by comparing the conventional band dispersions, 3D band dispersions, isoenergetic contours, and projected spin textures obtained from theoretical or experimental techniques with that obtained from the relevant model(s) [11, 10]. Below, we write a comprehensive model Hamiltonian for Rashba-Dresselhaus interactions in the first and third order and evaluate the eigenvalues and eigenstates corresponding to some useful combinations.

### v2

# Mapping Rashba and Dresselhaus spin-orbit interactions to inversion asymmetry in perovskite oxide heterostructures

@@ Abstract
Inversion asymmetry, combined with spin-orbit interaction, leads to Rashba or Dresselhaus effects or combinations that are promising for technologies based on antiferromagnetic spintronics. Since understanding the exact nature of spin-orbit interaction is crucial for developing a technology based on it, mapping the nature of inversion asymmetry with the type of spin-orbit interaction becomes the key. We simulate a perovskite oxide heterostructure LaAlOSrIrOSrTiO3 preserving the inversion symmetry within density functional theory to demonstrate the relation between the nature of inversion asymmetry and the corresponding Rashba or Dresselhaus-type interaction. With progressive distortion in the heterostructure, we find how the structure inversion asymmetry sets in with distorted bond lengths and angles, leading to the Rashba effect in the system in the presence of a microscopic electric field. Further, the introduction of tilted IrO6 octahedra leads to bulk inversion asymmetry, helping a combined Rashba-Dresselhaus interaction to set in. A comparison of the spin textures obtained from our DFT calculations and theoretical modeling helps us identify the exact nature of the interactions. In the context of perovskite heterostructures, our work demonstrates what type of distortions lead to bulk inversion asymmetry and what other distortions can lead only to structure inversion asymmetry by mapping different types of spin-orbit interactions to different distortions. Additionally, our work emphasizes the importance of spin texture in identifying the nature of spin-orbit interaction. It may serve as a guide to identifying different types of Rashba-like spin-orbit interactions.
@@ I Introduction
The surfaces and interfaces of perovskite oxides with the general formula ABO3 host Rashba-like interactions of various nature, order, and magnitude [10, 11, 12, 13, 14, 15]. Additionally, owing to the two-dimensional conducting system at the interface and matured synthesis techniques, electronics using perovskite oxide heterostructures emerged as a promising research direction [16, 17]. Interesting magnetic properties have also been observed at such interfaces, shaping them as potential candidates for technology [18, 19, 11]. A host of possible distortions in perovskite oxides make them an ideal subject for studying several interesting physical properties and their interrelation with distortion. For example, KTaO3 is usually found in a perfect cubic structure, YMnO3 shows Jahn-Teller (JT) distortion, SrTiO3 exhibits tetragonal distortion, and CaTiO3 hosts tilted TiO6 octahedra. Additionally, inversion asymmetry arises in the structure due to the formation of interfaces or surfaces. Microscopic electric fields developed in the layered polar oxides also help manifest the Rashba interaction. Thus, perovskite oxide heterostructures offer a platform to systematically investigate and verify the specific types of asymmetry and their implications on Rashba-like spin-orbit interaction.
@@ II Method
The work has been carried out using ab initio calculations based on density functional theory (DFT) and analytical model calculations, with the latter being described in detail in the appendix A. The DFT calculations for the LaAlOSrIrOSrTiO3 heterostructures considered here were performed using the projector augmented wave (PAW) method [20] in conjunction with a plane-wave basis set, as implemented in the vasp code [21, 22]. A generalized gradient approximation (GGA) due to Perdew et al. [23] along with a Hubbard- correction for the transition metal -orbitals due to Dudarev et al. [24], henceforth referred to as the GGA+ method, was invoked for approximating the exchange-correlation functional, with a moderate value of , 3, and 10 eV for Ir-, Ti-, and La- orbitals, respectively [11]. The Hubbard correction used for different elements serve different purposes. The empty La- states place themselves at a much lower energy than expected. We use a large value for these states to correct their placement in energy [19]. The relatively small value for the Ti- states helps correct the description of any transferred electron to those states [19]. Ir- states host a weak magnetism in the heterostructure when described within a small Hubbard- correction [11]. In order to make the results comparable with our previous works, we use the same values in our calculations for the heterostructure. Some key results were tallied against local density approximation (LDA) [25, 26] + Hubbard method, hereafter referred to as the LDA+ method. Unless mentioned otherwise for some specific cases, the lattice vectors and the atomic positions were optimized to minimize the stress and the Hellman-Feynman force on each atom to a tolerance of eV/Å, respectively, subject to a fixed lattice constant of 3.905 Å in the -plane to match that of a thick STO substrate in practice. -centered and -point meshes were used for the integrations over the Brillouin zone using improved tetrahedron method [27] for the calculations of three-dimensional (3D) band dispersions, isoenergetic contours and spin textures with and cells, respectively. Calculations accounting for spin-orbit interaction were performed with an energy convergence threshold of eV.
@@ III Results and Discussions
At first, we examine the spin-unpolarized density of states (DoS) and band structures of the heterostructure, shown in Fig. 2, for the three different structural forms mentioned above. The DoS and the projected DoS in Fig. 2(a) reveal the conducting nature of the heterostructure with the states near the Fermi level comprising Ir- orbitals hybridized with O- orbitals. The Ti- orbitals are found above the Fermi level, with the and orbitals higher in energy than the orbitals, as seen from the band dispersion in Fig. 2(b) [19]. While the orbitals are degenerate in the bulk perovskites, confinement along the -direction in the heterostructure elevates the and bands in energy. From the electronic structure, we deduce that the nominal oxidation states for Ti and Ir are and , respectively. The structural form accommodating JT-like tetragonal distortion reveals no significant change in the electronics structure (see Fig. 2(c), Fig. 2(d)), except the and bands coming closer to the bands, owing to the elongation of bonds along the -direction, thereby reducing the effect of the confinement. The structure with optimally distorted bond lengths and bond angles exhibits some significant changes in the electronic structure, as seen from the DoS and the band dispersion in Fig. 2(e) and Fig. 2(f), respectively. We find the Ti- and the Ir- bands overlapping in energy above the Fermi level. The energy difference between the bands and the , bands further reduce upon accommodating the distortions, and some of the otherwise degenerate bands near the -point become non-degenerate due to reduced symmetry (see Fig. 2(f)).
@@
In order to understand the Rashba-like spin-orbit interaction in the system with progressive asymmetry, we plot the energy as functions of and (so-called 3D bands) corresponding to the bands crossing the Fermi energy for the heterostructures with all three types of distortions, and the corresponding isoenergetic contours and spin textures in Fig. 5. Fig. 5(a) displays the plot corresponding to the bands crossing the Fermi level for the undistorted heterostructure. Fig. 5(b) and Fig. 5(c) show the corresponding isoenergetic contours and the projected spin directions marked using arrows arising from two different SrIrO3 layers. All the 3D bands and the isoenergetic contours exhibited in Fig. 5 show a four-fold rotational symmetry, typical of and bands in the plane [10]. The projected spin vectors at a crystal momentum are almost collinear and do not show a helical pattern in this case. The essential characteristic of a spin texture corresponding to Rashba-like spin-orbit interaction, viz. for a particular band is also not observed here. Instead, the projected spins corresponding to a specific band exhibit behavior. Therefore, we conclude that the antiferromagnetic arrangement of spins leads to the opposite spins from different bands in the spin texture, while no Rashba-like interaction is observed in the undistorted heterostructure due to the symmetry being preserved. Owing to preserved overall inversion symmetry and centrosymmetry at the point group level, the heterostructure with JT-like tetragonal distortions also does not exhibit any Rashba-like interaction, as confirmed from the dispersion plot, the isoenergetic contours, and the spin textures shown in Fig. 5(d),(e),(f). The Rashba-like spin helix obeying the condition for a particular band is not observed here, although the spin vectors far away from the Brillouin zone center are not always collinear with those near the zone center.
In the case of the heterostructure with distorted bond lengths and bond angles, we find from the 3D bands in Fig. 5(g), and the isoenergetic contours in Fig. 5(h) and Fig. 5(i) arising from two different SrIrO3 layers that the bands crossing the Fermi level are almost degenerate at and lines, but have a large separation along the diagonals of the Brillouin zone. The projected spin textures corresponding to these bands, displayed in Fig. 5(h) and Fig. 5(i), reveal the spins orienting themselves tangentially to a circle, obeying the feature for both layers of SrIrO3. Such orientation of the spins confirms predominantly linear Rashba-like interaction in the system (see the appendix) owing to the compromised structure inversion symmetry due to the distortions and the microscopic electric field [11].
@@ IV Conclusion
The role of inversion asymmetry in spin-orbit interaction-driven spin splitting, particularly bulk inversion asymmetry and structure inversion asymmetry leading to Dresselhaus and Rashba splitting, respectively, has earlier been theoretically established [6, 5]. However, mapping the commonly occurring structural distortions in perovskite structures to bulk or structure inversion asymmetry is often challenging in perovskite oxide heterostructures due to their complex nature. Moreover, similar splitting patterns arising from different spin-orbit interactions result in further confusion. However, the proposed technological applications based on such spin splittings critically depend on the exact nature of the interaction, making understanding the nature of interaction imperative. Our study illustrates how inversion asymmetry in the presence of a microscopic electric field [11] can lead to Rashba and Dresselhaus spin-orbit interaction using an example of a perovskite oxide heterostructure. Our first-principles calculations within density functional theory, 3D band dispersion, and a comparison of the spin textures obtained from our DFT calculations and theoretical models help us connect the structural distortions with the Rashba-like spin-orbit interaction pattern observed in such systems. Starting from a heterostructure of LaAlOSrIrOSrTiO3 simulated in such a way that preserves the inversion symmetry, we allowed progressive distortions in the bond lengths and the bond angles, without the IrO6 octahedra being tilted. Our results reveal linear Rashba spin-orbit interaction for the distorted heterostructure, suggesting only structure inversion asymmetry was realized through the distortions, while the bulk inversion symmetry was intact. However, considering a heterostructure allowed for tilted IrO6 octahedra, leading to bulk and structure inversion asymmetry manifesting as a combined Rashba and Dresselhaus spin-orbit interaction. A comparison of the spin texture obtained from our DFT calculations for a heterostructure and a spin texture predicted from our previous model calculation are in excellent agreement [11], implying the model’s validation. We have also considered various combinations of Rashba and Dresselhaus interactions, compiling an elaborate reference for the spin-splitting and spin textures corresponding to such interactions. Besides illustrating the connection between structural asymmetry and the Rashba-like spin-orbit interactions through an example of perovskite oxide heterostructures, our work helps classify via the spin textures similar spin-splitting due to spin-orbit interaction. Our results indicate that Jahn-Teller distortion or other distortions of bond length and bond angles in perovskite oxides do not lead to bulk inversion asymmetry, although it may be enough to break the overall inversion symmetry in a heterostructure. However, octahedral tilts observed in this perovskite heterostructure is a systematic structural distortion leading to bulk inversion asymmetry and Dresselhaus interaction in the heterostructure. Thus, we observe Rashba and Dresselhaus interactions coexisting in the heterostructure. Our techniques of analyzing the spin textures may help study topological materials and understand the corresponding transport properties.
@@ Appendix A A manual for identifying the nature of Rashba-like interactions
The spin-orbit interaction of the Rashba or Dresselhaus type arises due to broken inversion symmetry in the structure. The interaction Hamiltonian consists of products of the crystal momentum components and Pauli matrices that remain invariant under all the symmetry operations of the system, leading to spin-split bands and a helical spin texture. Therefore, one must first find the underlying symmetry and point group to find an appropriate spin Hamiltonian for a given system. Since spin splitting happens due to inversion asymmetry, the odd terms in , particularly the linear and the cubic terms, contribute the most. However, due to their structural complexity, finding the underlying symmetry may often be difficult in distorted perovskite oxide compounds, heterostructures, and superlattices. Moreover, the structural symmetry often does not predict the relative strength between the linear and the cubic terms. To avoid such confusion, here we collect some linear and cubic Rashba and Dresselhaus interaction terms and show the corresponding band dispersions, isoenergetic contours, and spin textures for the purpose of identifying the nature of the interaction in a given system.
Various Rashba-like spin-orbit interactions of different orders and their combinations may be identified by comparing the conventional band dispersions, 3D band dispersions, isoenergetic contours, and projected spin textures obtained from theoretical or experimental techniques with that obtained from the relevant model(s) [11, 10]. Below, we write a representative model Hamiltonian for Rashba-Dresselhaus interactions in the first and third order and evaluate the eigenvalues and eigenstates corresponding to some useful combinations. Since we consider a representative model Hamiltonian for illustration purposes, all the terms included in the model may not remain invariant under a given symmetry point group.